OpenStudy (anonymous):

the veranda is covered with tiles(30cm times 30cm) in 5 black and4 white tiles,how many black tiles are used to cover the veranda floor if that pattern is continued

OpenStudy (anonymous):

what is the total area of the veranda?

OpenStudy (anonymous):

the area is 4,2m time 1m=4,2m

OpenStudy (anonymous):

total area = \[4.2\operatorname{m}^2\]?

OpenStudy (anonymous):

the total area is 4.2 \[m ^{2}\]

OpenStudy (anonymous):

do you have an answer? I get 25 but I am rounding down as the answer I got is not an integer...

OpenStudy (anonymous):

so the answer is 25 black tiles

OpenStudy (amistre64):

Veranda is covered with tiles (30cm times 30cm): is 30 by 30 the measurement of each tile?

OpenStudy (anonymous):

yes in cm

OpenStudy (amistre64):

good, a centimeter is 1/100 of a meter. You said that the enire veranda is 4.2 meters squared right?

OpenStudy (anonymous):

yes.

OpenStudy (amistre64):

.03 (.03) = .0009 m^2 right?

OpenStudy (anonymous):

yes because you converted the cm to m

OpenStudy (amistre64):

thats my error... it should be .3(.3) = .09 m^2

OpenStudy (amistre64):

yes, we need to convert centimeters to meters to be able to determine how many m^2 one tile covers

OpenStudy (anonymous):

you right it is 0.3 because 30/100 =0.3m

OpenStudy (amistre64):

.09x=4.2? x = 4.2/.09 = 420/9 which is 46 and 2/3 tiles all together

OpenStudy (amistre64):

I know that 7 times 6 is equal to 42; so I assume that this veranda is NOT a square. and that it is either .6 meters by 7meters, or .7meters by 6 meters. Would that be a fair assumption to you?

OpenStudy (anonymous):

the veranda lenght is 4,2m by breadth 1m

OpenStudy (amistre64):

4.2 by 1 meter? thats good too :) either way we need 46/23 tiles all together.

OpenStudy (anonymous):

@amistre, is area under any curve found out by integrating the function?

OpenStudy (amistre64):

@think; yes. When we integrate a function, we are adding up all the little slices that it is made of. Which would equal its area.

OpenStudy (amistre64):

I get 3 and 1/3 tile in 1 meter.

OpenStudy (anonymous):

okay, so integrating=finding the area & differentiating= ?

OpenStudy (amistre64):

differentiating is ....finding the rate of change with respect to another variable. Differentiate with respect to (x) or (time). And thats finding the derivative.

OpenStudy (amistre64):

and I get 14 tiles to fit in 4.2 meters

OpenStudy (amistre64):

14(.3) = 4.2 right?

OpenStudy (anonymous):

30cm * 30cm=900cm 900cm divide by 100m =9m*4,2m= 40 tiles

OpenStudy (amistre64):

cant do it like that; you have to realize that the area changes and you no longer deal with 900 cm.. each tile is .3 meters wide; (.3)(.3) = .09 m^2 you have to use this measure of area. otherwise you mess yourself up. Does that make sense?

OpenStudy (amistre64):

900cm^2 is not easily converted to m^2 so convert it before you find the area.

OpenStudy (anonymous):

oh okay so the answer u gave me the 1st one is right

OpenStudy (amistre64):

Yes. .3 meters wide by .3 meters deep. .3x = 4.2 x = 14 tiles along the 4.2 meter side. .3x = 1 x = 1/.3 = 10/3 x = 3 and 1/3 tile to span 1 meter side. Do you agree with those numbers?

OpenStudy (anonymous):

yep,thanks i'll tel you the feed back tomrw ,pls help me with the last two parts of my homework

OpenStudy (amistre64):

black = 22 and white = 21 and 1/3 by those measurements

OpenStudy (amistre64):

whats the last two parts of your homework?

OpenStudy (amistre64):

Doh, I drew 13 tiles...let me reccount that..

OpenStudy (anonymous):

its going on grt though...

OpenStudy (amistre64):

You need 23 and 1/3 black tiles and 23 and 1/3 white tiles. Better buy 24 :)

OpenStudy (anonymous):

24 black tiles.right ...the front of the roof is in the shape of a isosceles triangle.the side of the roof are 50cm longer at the ends of the roof hiegh is 1,7m .calculate the lenght of the sides of the of the triangle

OpenStudy (amistre64):

is the height from ground level to the peak or is it from the ceiling line to the peak? This information only makes sense if we have a triangle that has a height of 1.7m and has little (50cm) overhangs extending beyond the base. Is the right?

OpenStudy (anonymous):

from the ceiling line to the peak.thats right the triangle height is 1,7m yes you right

OpenStudy (amistre64):

An isoTriangle has 60 degree corners; and the the sin(60) = 170cm/x(cm) x = 170/sin(60); then add the 50cm to get: x = 170/sin(60) + 50. Does that work for you?

OpenStudy (anonymous):

explain more

OpenStudy (amistre64):

246.3 cm is the answer I get. But to explain more: Have you learned about sine, cosine, and the functions of angles yet?

OpenStudy (amistre64):

or just the pythagorean theorum?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

Do you know the pythagorean theorum then? x^2 + y^2 = r^2 ?

OpenStudy (anonymous):

@amistre, do u know DeMorgan's laws?

OpenStudy (anonymous):

i dont know them

OpenStudy (amistre64):

@think; I have vaguely seen it, but I havent gone over it any. Whats it about?

OpenStudy (anonymous):

its related to sets & functions

OpenStudy (anonymous):

dont we start by converting the 50cm to meters

OpenStudy (amistre64):

Kab: in order to solve this problem; you need to know how to solve right triangles. and you solve them with a formula that says the height^2 + base^2 is equal to the slanted part^2

OpenStudy (amistre64):

@think; not in my repetoire yet :)

OpenStudy (amistre64):

and a right triangle is just half of a square. do you know that?

OpenStudy (anonymous):

bye guys..u both cont..gotta go

OpenStudy (amistre64):

Ciao think :)

OpenStudy (anonymous):

all d best wid ur discussion c u 'll the nxt time

OpenStudy (anonymous):

(H)^2+(B)^2 (1,7)^2+(0,05)^2 2.89m+0.0025m=2.89m^2 i converted the 50cm to meter which is 0,05

OpenStudy (amistre64):

you are on the right track, but I have to point something out to you first. for starters the 50 sm is something we will add at the end because it is not a part of the triangle to begin with. second; we dont know the B yet in order to find the length of the side of the roof. But we do know (because it is an isosolec(sp?) triangle) that the length of the roof is twice B. So lets set up out formula like this: (1.7)^2 + B^2 = (2B)^2 1.89 = 4B^2 - B^2 1.89 = 3B^2 1.89/3 = B^2 ...now square root both sides. sqrt(1.89/3) = B the length of the roof is twiceB 2(sqrt(1.89/3)) = 2B = length of one side of the roof; after you get this number add the overhang of (50cm.) Make sense?

OpenStudy (anonymous):

to make things easy for me may you please give the the hole sum and answer

OpenStudy (amistre64):

Yes, If I did it correctly; I get 2(170)/sqrt(3) + 50 196.3 cm + 50 cm = 246.3 cm is the roofs length on each side. or 2.463 meters

OpenStudy (anonymous):

do you stil have time for the last one, calculate the lenght and breadth of the parallelogram

OpenStudy (amistre64):

I got lotsa time... at least 4 hours :)

OpenStudy (amistre64):

lets go over to that posting for it tho.